OpenStudy (lgbasallote):

\[\large (x+y)dx + (x-y)dy = 0\] is the answer \[\large \frac{x^2}{2} + xy - \frac{y^2}{2} + c?\] im a little skeptic abt it :/

OpenStudy (unklerhaukus):

\[\frac{\partial}{\partial y}(x+y)=1\]\[\frac{\partial}{\partial x}(x-y)=1\]the equation is exact

OpenStudy (lgbasallote):

yes it is exact

OpenStudy (unklerhaukus):

\[f(x,y)=\int(x-y)\text dy+g(x)\]\[\qquad =xy-\frac{y^2}{2}+g(x)\] \[\frac{\partial f(x,y)}{\partial x}=y+g'(x)=(x+y)\] \[g'(x)=x\]\[g(x)=\frac{x^2}2\] \[f(x,y)=xy-\frac {y^2}2+\frac {x^2}2=c_1\] \[f(x,y)=\frac {x^2}2 +xy-\frac {y^2}2+c=0\]

OpenStudy (unklerhaukus):

i hope my method is clear enough

OpenStudy (lgbasallote):

it is...thanks...

OpenStudy (unklerhaukus):

yah, i only just learnt how to do these things myself

OpenStudy (lgbasallote):

me too *brofist*

OpenStudy (unklerhaukus):

indeed